7.1.37 Question Help A security expert claims that more than 15% of all homeowners have a home security alarm. State Ho and H7.1.38 Question Help A baseball team claims that the mean length of its games is at most 1.2 hours. State Ho and H, in words7.2.21-T 3 Question Help Find the critical value(s) and rejection region(s) for the type of z-test with level of significanceTest the claim about the population mean, H, at the given level of significance using the given sample statistics. Claim; u =A random sample of 79 eighth grade students scores on a national mathematics assessment test has a mean score of 294. This tA company claims that the mean monthly residential electricity consumption in a certain region is more than 880 kilowatt-hour
Answer #1

1) Answer C. The null hypothesis expressed in words is, "the proportion of all homeowners who own a home security is at most 0.15". The null hypothesis is expressed symbolically as "Ho : p ≤ 0.15."

--------

2) Answer B. The null hypothesis expressed in words is, "the mean length of a baseball team's game is at most 1.2 hours". The null hypothesis is expressed symbolically as "Ho : µ ≤ 1.2."

--------

3) Right tailed critical value, z crit = ABS(NORM.S.INV(0.02)) = 2.05

--------

4) x̅ = 48.6, σ = 3.43, n = 58

Null and Alternative hypothesis: Answer C.

Ho : µ = 50

H1 : µ ≠ 50

Critical value :

Two tailed critical value, z crit = ABS(NORM.S.INV(0.05/2)) = +/- 1.960

Reject Ho if z < -1.96 or if z > 1.96

Test statistic:

z = (x̅- µ)/(σ/√n) = (48.6 - 50)/(3.43/√58) = -3.1085

p-value :

p-value = 2*(1-NORM.S.DIST(ABS(-3.1085, 1) = 0.0019

Decision:

p-value < α, Reject the null hypothesis

---------

5) x̅ = 294, σ = 40, n = 79

Null and Alternative hypothesis: Answer C.

Ho : µ ≤ 285

H1 : µ > 285 (claim)

Critical value :

Right tailed critical value, z crit = ABS(NORM.S.INV(0.06) = 1.555

Reject Ho if z > 1.555

Test statistic:

z = (x̅- µ)/(σ/√n) = (294 - 285)/(40/√79) = 1.9998

p-value :

p-value = 1- NORM.S.DIST(1.9998, 1) = 0.0228

Decision:

p-value < α, Reject the null hypothesis

--------

6) x̅ = 900, σ = 129, n = 60

Null and Alternative hypothesis: Answer F.

Ho : µ ≤ 880

H1 : µ > 880 (claim)

Critical value :

Right tailed critical value, z crit = ABS(NORM.S.INV(0.1) = 1.282

Reject Ho if z > 1.282

Test statistic:

z = (x̅- µ)/(σ/√n) = (900 - 880)/(129/√60) = 1.2009

p-value :

p-value = 1- NORM.S.DIST(1.2009, 1) = 0.1149

Decision:

p-value > α, Do not reject the null hypothesis